find the common roots of these equations.
z3+2z2+2z+1=0
z1985+z100+1=0
 
(a) w, w2 
(b) 1, w, w2
(c) -1, w, w2
(d) -w, -w2
 

Grade:12

1 Answers

uday kiran
6 years ago
the answer is option A.
W+W^2+W^3=0 and W^3=1 are the formulas required to solve this problem.
substitute W in equation 1,we get W^3+2W^2+2W+1=0
which implies 1+2(W^2+W)+1=0
W^2+W=-W^3=-1
HENCE THIS BECOMES 2+2(-1)=0.hence W is root of given equation.similarly try with W^2.
now substitute W in equation 2.
the equation becomes W^1985+W^100+1=0
NOW DIVIDE 1985 BY 3.we get 2 as remainder.
divide 100 by 3.we get 1 as remainder.
now the equation becomes W^3*661(W^2)+W^3*33(W)+1=0
AS WE KNOW THAT W^3=1
now the equation becomes W^2+W+1=0
Hence proved.similarly try with W^2